# Microsoft cloud-based digitization workflow with rich metadata acquisition for cultural heritage objects

### v1

@@ Abstract
In response to several cultural heritage initiatives at the Jagiellonian University, we have developed a new digitization workflow in collaboration with the Jagiellonian Library (JL). The solution is based on easy-to-access technological solutions—Microsoft 365 cloud with MS Excel files as metadata acquisition interfaces, Office Script for validation, and MS Sharepoint for storage—that allows metadata acquisition by domain experts (philologists, historians, philosophers, librarians, archivists, curators, etc.) regardless of their experience with information systems. The ultimate goal is to create a knowledge graph that describes the analyzed holdings, linked to general knowledge bases, as well as to other cultural heritage collections, so careful attention is paid to the high accuracy of metadata and proper links to external sources. The workflow has already been evaluated in two pilots in the DiHeLib project focused on digitizing the so-called “Berlin Collection” and in two workshops with international guests, which allowed for its refinement and confirmation of its correctness and usability for JL. As the proposed workflow does not interfere with existing systems or domain guidelines regarding digitization and basic metadata collection in a given institution (e.g., file type, image quality, use of Dublin Core/MARC-21), but extends them in order to enable rich metadata collection, not previously possible, we believe that it could be of interest to all GLAMs (galleries, libraries, archives, and museums).
@@
Cultural heritage, Metadata, Knowledge graphs, GLAM, Digitization, Spreadsheets, Microsoft 365
@@ 1 Introduction and motivation
Jagiellonian University (JU), the oldest Polish university and one of the oldest universities in continuous operation in the world, holds in its collections a number of cultural heritage objects, including medieval manuscripts, social life documents representing a wide range of historical periods, memorabilia of former employees and students of the university, and a collection of authentic relics from the circle of ancient Mediterranean cultures. These collections have attracted greater research interest recently due to large cultural heritage initiatives at the JU, particularly the so-called flagship projects (FPs) [^1] See https://id.uj.edu.pl/en_GB/projekty-flagowe for more details. aimed at interdisciplinary and international cooperation to enrich the JU’s research ecosystem. Five of the 21 FPs are focused on cultural heritage research, including digitization, standardization of descriptions, integration of collections, and more detailed studies of selected items (cf. Sect. 3).
@@
With this in mind, we decided to take advantage of knowledge graphs [10] and ground the entire solution in the Semantic Web technologies [9] and the linked data approach [8, 12]. Adopting this approach will allow for the explicit definition of the meaning of individual metadata, the preservation of a flexible data model that can be easily adapted to the needs of individual collections or research objectives, the efficient integration of data across disciplines and organizations, and will facilitate future data processing in automated systems [15].
@@
During the development of metadata acquisition tools, a number of challenges and requirements had to be addressed, among which the most important were the tight time frame for tool development, the lack of resources to develop dedicated software, the need to use tools that are as familiar and accessible as possible to domain experts, and the scalability of the solution to collections with hundreds of thousands of objects. Taking into account all of these, we decided to base the solution on Microsoft 365 cloud and use Microsoft Excel spreadsheets as an interface for metadata collection, Office Script to facilitate the use of the spreadsheets, Python scripts for initial metadata validation, and dedicated space in Microsoft Sharepoint for data storage. The whole solution is complemented by a dedicated workflow based on the developed tools.
@@
The original contribution of the paper is the idea of a digitization workflow, based on easy-to-access technological solutions (Microsoft 365 cloud), that allows rich metadata acquisition by domain experts regardless of their experience with information systems. The proposed solution does not interfere with existing systems or domain guidelines regarding digitization and basic metadata collection in a given institution (e.g., file type, image quality, use of Dublin Core/MARC-21), but extends them in order to enable rich metadata collection, not previously possible. Therefore, we believe that it could be of interest to all GLAMs (galleries, libraries, archives, and museums). While the use of Excel for metadata collection is quite common, basing the entire digitization workflow on Microsoft 365 cloud to the best of our knowledge has not been described before.
@@
The remainder of the paper is structured as follows. The related work is summarized in Sect. 2. Sect. 3 introduces cultural heritage collections and digitization projects at Jagiellonian University. The requirements for the entire workflow are listed in Sect. 4. The tools used in the proposed solution are outlined in Sect. 5, while the actual digitization workflow is described in Sect. 6. Sect. 7 concludes the paper.
2 Related work
Harnessing the power of Semantic Web technology to integrate disparate cultural heritage (knowledge) bases requires well-thought-out shared vocabularies or ontologies. These shared resources establish common terminologies and relationships, facilitating interoperability and enhancing the effectiveness of data exchange and comprehension among disparate institutions. The two basic vocabularies are MARC-21, traditionally used in cataloging systems, and Dublin Core, traditionally used in digital libraries. However, they only allow for a basic description including title, author, year and place of creation, or access rights. More detailed metadata requires additional vocabularies or ontologies. Therefore, similar projects should be evaluated to determine whether they use vocabularies that can be adapted to the needs of the institution or collection under question.
One such project is the OntoBelliniLetters ontology of the Belliniano Civic Museum of Catania. Drawing on the main experience in this project, Cristofaro and Spampinato [3] report on the main concepts and relations used to develop the ontology “concerning the semantic organization of the corpus of Vincenzo Bellini’s correspondence letters” (p. 192). This ontology is inspired by CIDOC CRM, a conceptual framework and vocabulary for describing the relationships between cultural heritage objects, events, actors, and their contexts in a structured and interoperable way.
Another project worth mentioning is the attempt at digitization of cultural heritage resources in the German state of Brandenburg. Preuss [21] reports on some challenges of such an undertaking, including the variety of institutions containing relevant collections: 80 archives, 140 libraries, and 150 museums. A significant goal of such a project is the integration with European-wide portals, including Europeana and Kalliope. Cooperation plays a crucial role in addressing these challenges, allowing the pooling of resources and expertise across these diverse institutions. Through cooperative efforts, Brandenburg aims not only to digitize its cultural heritage, but also to seamlessly integrate it into larger European frameworks, fostering greater accessibility and appreciation of its rich historical and cultural legacy on a global scale. Both the targeted portals and the cooperative nature are inspirations for the projects undertaken currently at the Jagiellonian University.
The presentation of the related project already points out to target platforms for integration and shared vocabularies that would allow such integration. An instance of a standard for metadata representation is CIDOC CRM. The widespread use of CIDOC CRM and its foundational role in the development of ontologies for specific use cases within given institutions make it an essential conceptual framework to consider while developing any cultural heritage-related ontology. The fundamental category in CIDOC CRM is event. This event-centric approach offers expressiveness by capturing the cultural context through human actions, flexibility by allowing for dynamic and detailed records, and precision by enabling the representation of historical processes as documentable units, simplifying time processing and enhancing visualization (cf. [7]). Fig. 1 shows a graphical example of the representation of a sculpture according to CIDOC CRM.
Figure 1: Example of a sculpture modeling in CIDOC-CRM [2].
A target platform for integration with the data contained in the Jagiellonian Library is Kalliope. Kalliope is an online catalog for personal papers, manuscripts, and publishers’ archives that integrates metadata from institutions all around the German-speaking countries. Standardized data from partner institutions can be made available for retrieval in Kalliope given compliance with the Encoded Archival Description (EAD) format. The EAD is an international XML-based standard for the description and exchange of archival data. EAD possesses a hierarchical structure and makes it possible to describe the relation between wholes and parts in a collection. EAD contains elements for persons, families, places, and corporate bodies, these elements allow references to unique identifiers for such individuals. Kalliope utilizes links to the Integrated Authority File (Gemeinsame Normdatei, GND) of the German National Library to identify individuals. Fig. 2 shows an example of data encoded according to the EAD schema.
Figure 2: EAD example as presented in the official Kalliope documentation (adapted from https://kalliope-verbund.info/files/6560c72ae631ecfac4c60d06dde2f57a1b57ceb3.xml).
Finally, another portal with which collaboration is desired is Europeana. It is a digital cultural platform that integrates resources from more than 3000 institutions concerning European cultural heritage. Europeana’s metadata representation is based on the native Europeana Data Model (EDM). Three fundamental distinctive classes in EDM are the provided cultural heritage object (represented as edm:ProvidedCHO), the digital representation of the object (edm:WebResource), and the aggregation as a representation of the link between the former and the latter (ore:Aggregation). An example of an EDM representation of the Mona Lisa painting is presented in Fig. 3.
Figure 3: Provider aggregation with descriptive metadata in Europeana Data Model [13].
Besides inspiration on vocabulary as linking targets, similar digital cultural heritage projects also show tools for the entry of metadata by domain experts. Kim, Nakamura, and Watanave [17] present two systems for managing metadata in projects related to digital archives: Archivematica and Omeka S. Archivematica is an open-source digital preservation system designed to automate the process of ingesting, preserving, and providing access to digital archives. It is specifically tailored for the long-term preservation of digital materials, including documents, images, audio, and video files, among others. Omeka S specializes in providing a user-friendly platform for creating and curating digital exhibits and collections. One of the key features of Omeka S is its flexibility and extensibility, allowing users to customize metadata schemas and data entry forms to suit the unique needs of their projects. Furthermore, Omeka S facilitates collaborative metadata entry by supporting multi-user access and workflow management features. This enables domain experts and curatorial teams to collaborate effectively in cataloging and annotating digital materials, ensuring accuracy and consistency in metadata representation.
Despite the existence of such specialized software for the management of cultural heritage data, a popular option remains the use of more general Microsoft Excel. Jan [14] reports on the use of Excel in a case study of the investigation and management of community heritage resources. Spreadsheets were used to input details of heritage resources, encompassing site identifiers, titles, spatial representations, city identifiers, and descriptions. The collected metadata attempted to follow some interoperable schemas, including CIDOC CRM.
Utilizing Excel spreadsheets for metadata collection offers the benefit of a user-friendly interface, reducing the learning curve for participants. This advantage is particularly valuable in projects lacking designated domain experts, where individuals are enlisted on an ad hoc basis for specific tasks as they arise during project execution. In such scenarios, the familiarity of Excel’s interface enhances efficiency by enabling quick adaptation and seamless participation, even among those with varied levels of technical expertise. One way shown by Jan [14] to ensure and enhance the quality of data entered in spreadsheets is by using Python scripts to extract, validate and integrate it with other resources such as ontologies, or another type of shared vocabularies.
Finally, Kovalenko, Serral and Biffl’s [19] discussion of tools for ontology population using spreadsheet data already shows some evaluated alternatives to the transformation of the collected metadata. In a more recent approach, Denisova, Dorodnykh and Yurin [4] present the OntoGen tool for creating ontologies by transforming spreadsheet data. Another tool worth mentioning is Owlready, a Python package that provides convenient manipulation and population of OWL ontologies (cf. [20]).
This brief overview of related projects begins to outline a workflow for managing digital cultural heritage at Jagiellonian University: domain experts gather metadata across various fields and input it into Excel spreadsheets. Python scripts are then employed to both validate the metadata and populate an ontology. This ontology is designed to be interoperable with Kalliope, Europeana, and CIDOC CRM. Before delving into the detailed explanation of this workflow in Sect. 4-LABEL:sec:evaluation, it is valuable to examine the current cultural heritage landscape at Jagiellonian University.
@@ 3 Cultural heritage collections and digitization landscape at Jagiellonian University
From the point of view of cultural heritage collections, the Jagiellonian Library (JL) should be considered a central part of the Jagiellonian University. Although it was not founded together with the University in 1364, the Library is a successor to various university libraries and collections that have existed since the 15th century. For many years, it also played the role of a national library, the so-called bibliotheca patria [1]. Now, JL holds 6,564,628 items [^2] For official numbers, see: https://bj.uj.edu.pl/en_GB/about-the-library/mission-history-and-collections/collections/jagiellonian-university-library-system-holdings-in-numbers. in its collections (not including items stored in faculty libraries), with many cultural heritage treasures including 11th century documents, collection of incunabula or Dutch, Italian and French engravings [1].
One of the most notable collections in the holdings of the Jagiellonian Library is the so-called “Berlin Collection” (affectionately termed “Berlinka” in Polish). It is a collection of over 500,000 historical documents that, before the end of World War II, belonged to the Prussian State Library (now State Library of Berlin) but is currently held by the Jagiellonian Library (for a detailed history of the collection, see [16]). The Autograph Collection, a premier segment of the Berlin Collection, is an assembly of unique manuscript materials from the 15th to the 20th century. It comprises approximately 300,000 autographs, including letters from Martin Luther, Johann Wolfgang von Goethe, Georg Wilhelm Friedrich Hegel, Heinrich von Kleist, Jacob and Wilhelm Grimm, Alexander and Wilhelm von Humboldt, Marie Antoinette, and many other poets, scientists, royals, statesmen, and philosophers. The collection not only exhibits the activity of particular authors, but also gives insight into the intellectual exchanges and development of cultural and ideological movements in Germany and in the whole Europe. The documents within the collection were already organized: they were divided into units associated with individual persons. Unit shelfmarks, assigned by Berlin librarians, are the names of these individuals preceded by the designation “SA,” (Sammlung Autographa) for example, “SA, Keppler, Johannes”. There are about 30,000 units within the Autograph Collection; the sizes of individual units vary greatly; some are extensive and consist of hundreds of documents, while others contain only a few, sometimes even just one.
@@
Jagiellonian University’s heritage collections are also spread in many places outside the Jagiellonian Library. For example, documents and photos of university staff and students are kept in the JU Archive, the JU Museum keeps a collection of memorabilia, the mediterranean and prehistoric archaeological collections are kept at the JU Institute of Archaeology [25]. There are also the photo collection of the Institute of Art History, the Ignacy Jan Paderewski Center for the Documentation of 19th- and 20th-century Polish Music, and many other collections [26].
@@
Most of the aforementioned cultural heritage objects have not yet been digitized. To illustrate, JL has digitized and made available to the public 830,887 objects. Considering the special collections alone—that is, excluding collections of books and periodicals, many of which cannot be digitized due to legal restrictions—comprising 2,296,930 objects, this means that more than 60 percent of the collections are still not digitized.
@@
These collections have attracted greater research interest recently due to large cultural heritage initiatives at the JU, particularly the so-called flagship projects (FPs) [^3] See https://id.uj.edu.pl/en_GB/projekty-flagowe for more details. aimed at interdisciplinary and international cooperation to enrich the JU’s research ecosystem. Five of the 21 FPs are focused on cultural heritage research, including digitization, standardization of descriptions, integration of collections, and more detailed studies of selected items (cf. Sect. 3).
However, it should be emphasized that significant initiatives for cultural heritage are currently underway at the Jagiellonian University. The Jagiellonian Center for Digital Humanities has been established, and 5 of the 21 flagship projects (see Sect. 1) cover topics related to the digitization of cultural heritage objects:
@@
    Mare Nostrum Lab (https://mare.id.uj.edu.pl/en_GB/) is aimed at the research related to the cultural heritage of Mediterranean societies.
  2. 2.
    Digital Humanities Lab (https://dhlab.id.uj.edu.pl/en_GB/) goal is to provide a stable and standardized digital editing platform for all JU scholars and students.
  3. 3.
    European Heritage in the Jagiellonian Library: Digital Authoring of the Berlin Collections (DiHeLib) (https://dihelib.id.uj.edu.pl/en_GB/) aims at cataloging, digitizing, sharing, and conducting research on the Autograph Collection of the “Berlinka” [23].
  4. 4.
    Cultural Heritage Exploration and Retrieval with Intelligent Systems at Jagiellonian University (CHExRISH) will provide a solution for combining a variety of JU collections oriented around the university’s history in a unified interface with Semantic Web technologies.
  5. 5.
    Critical Heritage Studies Hub (CHSH) (https://crihestu.id.uj.edu.pl/en_GB/) is not aimed at digitization, but among its goals is the creation of thesauri in SKOS (Simple Knowledge Organization System) format for cultural heritage collections.
The presented cultural heritage landscape is accompanied by various systems used to store metadata about objects:
  1. 1.
    JL uses the Alma interlibrary catalog (https://katalogi.uj.edu.pl/), which stores data in MARC 21 format, and the Jagiellonian Digital Library (https://jbc.bj.uj.edu.pl/) based on the dLibra system, which stores PDFs and metadata in Dublin Core format.

### v2

# Cloud-based digitization workflow with rich metadata acquisition for cultural heritage objects

@@ Abstract
In response to several cultural heritage initiatives at the Jagiellonian University, we developed a new digitization workflow in collaboration with the Jagiellonian Library (JL). The solution is based on easy-to-access technological solutions—Microsoft 365 cloud with MS Excel files as metadata acquisition interfaces, Office Script for validation, and MS Sharepoint for storage—that allows metadata acquisition by domain experts regardless of their experience with information systems. The ultimate goal is to create a knowledge graph that describes the analyzed collections, linked to general knowledge bases, as well as to other cultural heritage collections, so careful attention is paid to the high accuracy of metadata and proper links to external sources. The workflow was evaluated in two pilot studies and in two workshops, which allowed for its refinement and confirmation of its correctness and usability for JL. The knowledge graph created as a result of these pilot studies was made available in a public git repository. As the proposed workflow does not interfere with existing systems or domain guidelines regarding digitization and basic metadata collection in a given institution, but extends them in order to enable rich metadata collection, not previously possible, we believe that it could be of interest to all GLAMs.
@@
Cultural heritage , Metadata , Knowledge graphs , GLAM , Digitization , Spreadsheets
journal: The Journal of Academic Librarianship
[…]
[1]organization=Department of Human-Centered Artificial Intelligence, Institute of Applied Computer Science, Faculty of Physics, Astronomy and Applied Computer Science, Jagiellonian University, addressline=S. Łojasiewicza 11, city=Kraków, postcode=30-348, country=Poland
[…]
[2]organization=Faculty of Humanities, AGH University of Krakow, addressline=Czarnowiejska 36, city=Kraków, postcode=30-054, country=Poland
@@ 1 Introduction and motivation
Jagiellonian University (JU), the oldest Polish university and one of the oldest universities in continuous operation in the world, holds in its collections a number of cultural heritage objects, including medieval manuscripts, social life documents representing a wide range of historical periods, memorabilia of former employees and students of the university, and a collection of authentic relics from the circle of ancient Mediterranean cultures. These collections have attracted greater research interest recently due to large cultural heritage initiatives at the JU, particularly the so-called flagship projects (FPs) [^1] See https://id.uj.edu.pl/en_GB/projekty-flagowe for more details. aimed at interdisciplinary and international cooperation to enrich the JU’s research ecosystem. Five of the 21 FPs are focused on cultural heritage research, including digitization, standardization of descriptions, integration of collections, and more detailed studies of selected items (cf. Sect. 2).
@@
With this in mind, we decided to take advantage of knowledge graphs (Hogan et al., 2021) and ground the entire solution in the Semantic Web technologies (Hitzler, 2021) and the linked data approach (Heath and Bizer, 2011; Hyvönen and Rantala, 2019). Adopting this approach will allow for the explicit definition of the meaning of individual metadata, the preservation of a flexible data model that can be easily adapted to the needs of individual collections or research objectives, the efficient integration of data across disciplines and organizations, and will facilitate future data processing in automated systems (Janowicz et al., 2015).
@@
During the development of metadata acquisition tools, a number of challenges and requirements had to be addressed, among which the most important were the tight time frame for tool development, the lack of resources to develop dedicated software, the need to use tools that are as familiar and accessible as possible to domain experts, and the scalability of the solution to collections with hundreds of thousands of objects. Taking into account all of these, we decided to base the solution on Microsoft 365 cloud and use Microsoft Excel spreadsheets as an interface for metadata collection, Office Script to facilitate the use of the spreadsheets, Python scripts for initial metadata validation, dedicated space in Microsoft Sharepoint for data storage, and Python scripts for ontology population from gathered data. The whole solution is complemented by a dedicated workflow based on the developed tools.
@@
The original contribution of the paper is the prototype of a digitization workflow, based on easy-to-access technological solutions (Microsoft 365 cloud), that allows rich metadata acquisition by domain experts regardless of their experience with information systems, resulting in the creation of knowledge base filled with gathered metadata. The proposed solution does not interfere with existing systems or domain guidelines regarding digitization and basic metadata collection in a given institution (e.g., file type, image quality, use of Dublin Core/MARC 21), but extends them in order to enable rich metadata collection, not previously possible. Therefore, we believe that it could be of interest to all GLAMs (galleries, libraries, archives, and museums). While the use of Excel for metadata collection is quite common, basing the entire digitization workflow on Microsoft 365 cloud to the best of our knowledge has not been described before.
@@
The remainder of the paper is structured as follows. Sect. 2 introduces selected cultural heritage collections and digitization projects at Jagiellonian University. The related work is summarized in Sect. 3. The requirements for the entire workflow are listed in Sect. 4. The tools used in the proposed solution are outlined in Sect. 5, while their evaluation during actual digitization workflow resulting in the first version of knowledge base is described in Sect. 6. Finally, Sect. 7 concludes the paper.
@@ 2 Cultural heritage collections and digitization landscape at Jagiellonian University
From the point of view of cultural heritage collections, the Jagiellonian Library (JL) should be considered a central part of the Jagiellonian University. Although it was not founded together with the University in 1364, the Library is a successor to various university libraries and collections that have existed since the 15th century. For many years, it also played the role of a national library, the so-called bibliotheca patria (Bakowska, 2005). Now, JL holds 6,798,156 items [^2] For official numbers, see: https://bj.uj.edu.pl/en_GB/about-the-library/mission-history-and-collections/collections/jagiellonian-university-library-system-holdings-in-numbers. in its collections (not including items stored in faculty libraries), with many cultural heritage treasures including 11th century documents, collection of incunabula or Dutch, Italian and French engravings (Bakowska, 2005).
@@
Jagiellonian University’s heritage collections are also spread in many places outside the Jagiellonian Library. For example, documents and photos of university staff and students are kept in the JU Archive, the JU Museum keeps a collection of memorabilia, the mediterranean and prehistoric archaeological collections are kept at the JU Institute of Archaeology (Woźny and Dzięgielewski, 2018). There are also the photo collection of the Institute of Art History, the Ignacy Jan Paderewski Center for the Documentation of 19th- and 20th-century Polish Music, and many other collections (Zięba, 2020).
@@
Most of the aforementioned cultural heritage objects have not yet been digitized. To illustrate, JL has digitized and made available to the public 904,624 objects. Considering the special collections alone—that is, excluding collections of books and periodicals, many of which cannot be digitized due to legal restrictions—comprising 2,390,834 objects, this means that more than 60 percent of the collections are still not digitized.
@@
These collections have attracted greater research interest recently due to large cultural heritage initiatives at the JU, particularly the so-called flagship projects (FPs) [^3] See https://id.uj.edu.pl/en_GB/projekty-flagowe for more details. aimed at interdisciplinary and international cooperation to enrich the JU’s research ecosystem. Five of the 21 FPs are focused on cultural heritage research, including digitization, standardization of descriptions, integration of collections, and more detailed studies of selected items. This cultural heritage landscape is accompanied by various systems used to store metadata about objects:
@@
    JL uses the Alma interlibrary catalog (https://katalogi.uj.edu.pl/), which stores data in MARC 21 format, and the Jagiellonian Digital Library (https://jbc.bj.uj.edu.pl/) based on the dLibra system, which stores scanned texts in PDF format along with their metadata in the Dublin Core standard.
@@
3 Related work
A common starting point for the decision of what software to include in the digitization workflow for metadata acquisition is evaluating existing digital library softwares (DLS). This evaluation consists of a mapping of institutional requirements to features offered by different tools. Among institutional requirements are the target format for metadata, the extent to which an institution is open for changes in their existing procedures and the training capabilities for introducing new interfaces for those performing the metadata entry operations. Among features of DLS are the supported metadata standards, licensing model, user interface design for metadata entry, controlled vocabularies, learning curve, and their capacity for future customization (Mukherjee and Das, 2020).
As discussed above, the limitations of the current digitization workflow of the Jagiellonian Library restricts the manual metadata entry process and storage to the use of tools present in the Microsoft 365 cloud package. Despite this restriction, an evaluation of existing DLS and a presentation of its current uses can still be useful for an outline of the target functionalities of the Microsoft cloud-based metadata acquisition workflow.
Verma and Kumar (2018) highlight three widespread DLS: DSpace [^4] https://www.dspace.org/, Greenstone [^5] http://www.greenstone.org/, and ePrints [^6] https://www.eprints.org/. DSpace is an open-source digital repository platform designed to capture, store, manage, and share digital content, widely used by academic institutions, libraries, and organizations for preserving scholarly outputs and institutional assets. It supports customizable metadata standards, interoperability protocols like OAI-PMH, and offers a user-friendly interface for managing collections. Verma and Kumar (2018) affirm that there are more than 1000 repositories which are using DSpace in at least one part of the development process. Chen et al. (2012) have reported on the process of using DSpace in the implementation of a Taiwan Digital Library History Library as both a metadata repository and a user-interface. Kurtz (2010) examines the influence of DSpace in The University of New Mexico, The University of Washington and Ohio State University as a metadata acquisition software and repository in the metadata quality of created records. From this brief investigation, DSpace appears to be a frequently chosen option due to its flexibility, scalability, and strong support for metadata management, including the use of linked data service and compatibility with discovery systems.
Greenstone is another open-source DLS that enables the creation, management, and dissemination of digital collections, and supports metadata formats including Dublin Core, as well as the OAI-PMH interoperability protocol. Greenstone provides a user interface for cataloging (i.e., inserting and editing records) and browsing digital collections, however it does not natively support linked data. A comprehensive overview of Greenstone current advancements can be found in (Bainbridge and Witten, 2020). A list of examples of libraries’ usage of Greenstone can be found in https://www.greenstone.org/examples. Finally, ePrints is also a open-source DLS for creating and managing digital repositories, and includes customizable metadata formats and interoperability, exposition of its metadata in RDF, integration with third-party services such as ORCID and a user-friendly interface and strong search capabilities for both end users and administrators. Pramudyo and Hendrawan (2020) present the use of ePrints in the Universitas Brawijaya Library as part of a process of metadata interoperability between several libraries using different digital library software.
An alternative software with a proprietary licensing model is CONTENTdm [^7] https://www.oclc.org/en/contentdm.html, a digital asset management software used by libraries to store, manage, and provide access to digitized collections that offers extensive support and hosting services for its users. Payant et al. (2020) consists of a report that portrays the metadata fields, mappings, and resources used for creating and managing metadata in Utah State University’s CONTENTdm digital collections. Bahnemann et al. (2021) share the findings of a pilot project investigating conversion methodologies for metadata originating from CONTENTdm into linked data.
The four aforementioned widespread DLS were developed and released in the late 1990s and early 2000s, that is, slightly before the publishing RDF 1.0 and OWL 1.0 specification. In the coming years, linked data became an important part of libraries’ endeavor for making their records more widely available, findable and interoperable (Hallo et al., 2016). While DSpace, ePrints and, to a lesser extent, CONTENTdm have provided updates for these tools supporting linked data solutions, other DLS have emerged with their functionalities based on the principles of linked data.
Omeka-S [^8] https://omeka.org/s/ is an open-source platform for managing and sharing digital collections, including their metadata. It supports Linked Open Data with features like customizable metadata schemas, RDF mapping, and external LOD system interoperability. It possesses a user interface that allows users to manage, create, and display exhibits. In Popović et al. (2020) one can read about the experience of implementing Omeka-S as data insertion, storage and interface at the Faculty of Mining and Geology Digital Repository at the University of Belgrade. Another example of usage of Omeka-S for semantic annotation, data storage and data browsing and querying can be found in Bikakis et al. (2021) for the Archives Henri Poincaré. Another LOD-based DLS is Arches-HIP [^9] https://arches-hip.readthedocs.io/en/latest/, an open-source platform for managing heritage data, emphasizing LOD principles for interoperability. It stands out due to its ability to integrate spatial analysis and visualization tools. In Yang et al. (2018) Arches-HIP is used for the registration and integration of Taiwan’s cultural heritage metadata into CIDOC CRM-based ontologies and presentation through a web platform in spatially meaningful representations. Finally, other tools used in in GLAM-related metadata tasks are, among others, CollectionSpace [^10] https://collectionspace.org/, Archivematica [^11] https://www.archivematica.org/ and OpenRefine [^12] https://openrefine.org/.
In some cases, the metadata acquisition workflow with the aforementioned DLS’ is complemented by using Microsoft tools, including MS Excel. Jan (2018) demonstrates the use of Excel for structuring and validating heritage resource data, such as site IDs, names, geometries, and descriptions, before integrating the data into advanced platforms like Arches-HIP for geospatial analysis and management. Moreover, Kuźma and Mościcka (2020) mention the use of Excel to process metadata of 35,092 topographic maps from the National Library of Poland. The precise role of Excel in this workflow, however, is not clear, as the authors only mentioned that “scripts and formulas were defined to classify and analyze obtained data” (Kuźma and Mościcka, 2020, p. 5). Besides these minor uses of MS Excel, no work has been found containing a thorough description of a metadata acquisition workflow with Microsoft 365 cloud package tools on the forefront.
Utilizing Excel spreadsheets for metadata collection offers the benefit of a user-friendly interface, reducing the learning curve for participants. This advantage is particularly valuable in projects lacking designated domain experts, where individuals are enlisted on an ad hoc basis for specific tasks as they arise during project execution. In such scenarios, the familiarity of Excel’s interface enhances efficiency by enabling quick adaptation and seamless participation, even among those with varied levels of technical expertise. One way shown by Jan (2018) to ensure and enhance the quality of data entered in spreadsheets is by using Python scripts to extract, validate and integrate it with other resources such as ontologies, or another type of shared vocabularies.
Finally, Kovalenko et al. (2013) discussion of tools for ontology population using spreadsheet data already shows some evaluated alternatives to the transformation of the collected metadata. In a more recent approach, Denisova et al. (2022) present the OntoGen tool for creating ontologies by transforming spreadsheet data. Another tool worth mentioning is Owlready, a Python package that provides convenient manipulation and population of OWL ontologies (cf. Lamy, 2017).
Concluding, the review of existing metadata acquisition softwares and workflows shows that the main functionalities present in such systems are the support customizable metadata standards, such as Dublin Core and RDF, including validation for different datatypes. They also provide user-friendly interfaces for manual and automated metadata entry. Furthermore, some systems incorporate Linked Data principles, enabling metadata enrichment. In the following sections we describe our solution for offering similar features on a Microsoft cloud-based environment.